# Piecewise linear constitutive relations for stretch-limited elastic strings

### v1

@@ Abstract.
This study proposes a simple and novel class of stretch-limiting constitutive relations for perfectly flexible elastic strings drawing from modern advances in constitutive theory for elastic bodies. We investigate strings governed by constitutive relations where stretch is a bounded, piecewise linear function of tension, extending beyond the traditional Cauchy elasticity framework. Our analysis includes explicit solutions for both catenaries and longitudinal, piecewise constant stretched motions.
@@ 1. Introduction
This paper focuses on the physically important but simpler setting in which classical and modern advances in constitutive theory for elastic bodies can be explored: perfectly flexible strings. Interest in the study of the mechanics of strings and fibers has a storied history; it attracted the attention of da Vinci, Galileo, Euler, the Bernoullis, Savart, Lagrange, Leibnitz, Ricatti, Huygens, Navier, Poisson, Lamé and other distinguished scientists. The interested reader can find a detailed account of the history of investigations into the mechanics of strings/fibers/wires in the work of Todhunter [34]. All of these early studies were within the context of special theories and not within the purview of general nonlinear theories of elasticity. Studies within the context of nonlinear theories are too numerous to mention and one can find copious references to both recent and older studies mentioned above in the treatise by Antman [1].
@@
In [29], the author established existence and multiplicity results for stretch-limited catenaries, strings suspended between two points, a problem that attracted the interests of the likes of Galileo, Bernoulli, Huygens, Leibniz, among others. He found that the multiplicity for such stretch-limited strings deviates significantly from the case of traditional classical elastic strings satisfying (1.1). The work [30] studied the propagation of longitudinal waves in a semi-infinite stretch-limited string with the finite end held fixed and the other end asymptotically in tension. There, he proved the orbital asymptotic stability of an explicit two-parameter family of piecewise constant stretched motions.
@@
In this work, we introduce a simple class of stretch-limiting, piecewise linear constitutive relations for elastic strings of the form $\nu=\hat{\nu}(N)$. The organization of the paper is as follows. In Section 2, we briefly review the kinematics, balance laws and constitutive theory governing perfectly flexible elastic strings. In Section 3, we present a simple form of $\hat{\nu}(N)-1$ that limits the amount of compression and extension the string can undergo, is linear in $N$ when the string is extensible, and has different tangent moduli (i.e., proportionality constants) for extension versus compression (see (3.2)). In Section 4, we set up simple boundary value problems that are solved in Section 5 (catenaries) and Section 6 (piecewise constant stretched motions) from the semi-inverse point of view.
@@ 4.1. Alternative forms of the field equations
In the case that $N<0$, the only modification we need to consider to the above equation is to use $\hat{\nu}(-|\boldsymbol{n}(\xi,t)|)$ instead $\hat{\nu}(|\boldsymbol{n}(\xi,t)|)$.
@@ 5. Catenaries
For each component ${r}_{i}$ of ${\boldsymbol{r}}$ (the components of ${\boldsymbol{n}}_{0}$ are ${n}_{0_{i}}$), we obtain
${r}_{1}({s})$ $=$ ${n}_{0_{1}}\ell({s}),\quad{r}_{3}({s})={n}_{0_{3}}\ell({s}),$ (5.6)
${r}_{2}({s})$ $=$ $\frac{1}{2\zeta}\left[2\sqrt{{n}_{0_{1}}^{2}+\left({n}_{0_{2}}+\zeta{s}\right)^{2}+{n}_{0_{3}}^{2}}+\zeta{s}\left(2{n}_{0_{2}}+\zeta{s}\right)-2\sqrt{{n}_{0_{1}}^{2}+{n}_{0_{2}}^{2}+{n}_{0_{3}}^{2}}\right],$ (5.7)
@@
$\ell({s})={s}+\frac{1}{\zeta}\ln\left[\frac{{n}_{0_{2}}+\zeta{s}+\sqrt{{n}_{0_{1}}^{2}+\left({n}_{0_{2}}+\zeta{s}\right)^{2}+{n}_{0_{3}}^{2}}}{{n}_{0_{2}}+\sqrt{{n}_{0_{1}}^{2}+{n}_{0_{2}}^{2}+{n}_{0_{3}}^{2}}}\right].$ (5.8)
@@
which can be used to find ${n}_{0_{1}}$, ${n}_{0_{2}}$ and ${n}_{0_{3}}$.
@@
In this section we study the case that there exist $\tilde{{s}}$ and $\breve{{s}}$ such that $10<\tilde{{s}}<1-\breve{{s}}<$ 1 and
@@
${r}_{1}({s})$ $=$ ${n}_{0_{1}}\ell_{A}({s}),\quad{r}_{3}({s})={n}_{0_{3}}\ell_{A}({s}),$ (5.14)
${r}_{2}({s})$ $=$ $\frac{\nu_{1}}{\zeta}\left[\sqrt{{n}_{0_{1}}^{2}+\left({n}_{0_{2}}+\zeta{s}\right)^{2}+{n}_{0_{3}}^{2}}-\sqrt{{n}_{0_{1}}^{2}+{n}_{0_{2}}^{2}+{n}_{0_{3}}^{2}}\right],$ (5.15)
@@
$\ell_{A}({s})=\frac{\nu_{1}}{\zeta}\ln\left[\frac{{n}_{0_{2}}+\zeta{s}+\sqrt{{n}_{0_{1}}^{2}+\left({n}_{0_{2}}+\zeta{s}\right)^{2}+{n}_{0_{3}}^{2}}}{{n}_{0_{2}}+\sqrt{{n}_{0_{1}}^{2}+{n}_{0_{2}}^{2}+{n}_{0_{3}}^{2}}}\right].$ (5.16)
@@
${r}_{1}({s})$ $=$ ${n}_{0_{1}}\ell_{B}({s})+{r}_{0_{1}},\quad{r}_{3}({s})={n}_{0_{3}}\ell_{B}({s})+{r}_{0_{3}},$ (5.18)
${r}_{2}({s})$ $=$ $\frac{1}{2\zeta}\left[2\sqrt{{n}_{0_{1}}^{2}+\left({n}_{0_{2}}+\zeta{s}\right)^{2}+{n}_{0_{3}}^{2}}+\zeta{s}\left(2{n}_{0_{2}}+\zeta{s}\right)\right.$ (5.19)
$\left.-2\sqrt{{n}_{0_{1}}^{2}+\left({n}_{0_{2}}+\zeta\tilde{{s}}\right)^{2}+{n}_{0_{3}}^{2}}-\zeta\tilde{{s}}\left(2{n}_{0_{2}}+\zeta\tilde{{s}}\right)\right]+{r}_{0_{2}},$
@@
$\ell_{B}({s})={s}-\tilde{{s}}+\frac{1}{\zeta}\ln\left[\frac{{n}_{0_{2}}+\zeta{s}+\sqrt{{n}_{0_{1}}^{2}+\left({n}_{0_{2}}+\zeta{s}\right)^{2}+{n}_{0_{3}}^{2}}}{{n}_{0_{2}}+\zeta\tilde{{s}}+\sqrt{{n}_{0_{1}}^{2}+\left({n}_{0_{2}}+\zeta\tilde{{s}}\right)^{2}+{n}_{0_{3}}^{2}}}\right].$ (5.20)
@@
${r}_{1}({s})$ $=$ ${n}_{0_{1}}\ell_{C}({s})+\hat{{r}}_{0_{1}},\quad{r}_{3}({s})={n}_{0_{3}}\ell_{C}({s})+\hat{{r}}_{0_{3}},$ (5.22)
${r}_{2}({s})$ $=$ $\frac{\nu_{1}}{\zeta}\left\{\sqrt{{n}_{0_{1}}^{2}+\left({n}_{0_{2}}+\zeta{s}\right)^{2}+{n}_{0_{3}}^{2}}-\sqrt{{n}_{0_{1}}^{2}+\left[{n}_{0_{2}}+\zeta\left(1-\breve{{s}}\right)\right]^{2}+{n}_{0_{3}}^{2}}\right\}+\hat{{r}}_{0_{2}},$ (5.23)
@@
$\ell_{C}({s})=\frac{\nu_{1}}{\zeta}\ln\left\{\frac{{n}_{0_{2}}+\zeta{s}+\sqrt{{n}_{0_{1}}^{2}+\left({n}_{0_{2}}+\zeta{s}\right)^{2}+{n}_{0_{3}}^{2}}}{{n}_{0_{2}}+\zeta{s}+\sqrt{{n}_{0_{1}}^{2}+\left[{n}_{0_{2}}+\zeta\left(1-\breve{{s}}\right)\right]^{2}+{n}_{0_{3}}^{2}}}\right\}.$ (5.24)
In the above results for ${r}_{i}({s})$ we have the following list of 11 unknown parameters: $\tilde{{s}}$, $\breve{{s}}$, ${r}_{0_{1}}$, ${r}_{0_{2}}$, ${r}_{0_{3}}$, $\hat{{r}}_{0_{1}}$, $\hat{{r}}_{0_{2}}$, $\hat{{r}}_{0_{3}}$, ${n}_{0_{1}}$, ${n}_{0_{2}}$ and ${n}_{0_{3}}$. These can be found from the 6 continuity conditions
@@
${n}_{0_{1}}^{2}+\left({n}_{0_{2}}+\zeta\tilde{{s}}\right)^{2}+{n}_{0_{3}}^{2}={N}_{1}^{2},$ (5.27)
${n}_{0_{1}}^{2}+\left[{n}_{0_{2}}+\zeta\left(1-\breve{{s}}\right)\right]^{2}+{n}_{0_{3}}^{2}={N}_{1}^{2}.$ (5.28)
From (5.25)-(5.28), we have 11 equations determining the list of 11 unknowns.
@@ 7. Final remarks
In the present work, we proposed a new type of piecewise linear constitutive relation for strings that limits the amount of extension and compression the string can undergo. Moreover, the tangent moduli relating tension to stretch are different for extension versus compression. We solved static and dynamic problems from the semi-inverse perspective with solutions to the latter having piecewise constant stretch. The full dynamic problem will be investigated in a future work, using a numerical method presented in [3].

### v2

@@ Abstract.
This study proposes a simple and novel class of stretch-limiting constitutive relations for perfectly flexible elastic strings drawn from modern advances in constitutive theory for elastic bodies. We investigate strings governed by constitutive relations where stretch is a bounded, piecewise linear function of tension, extending beyond the traditional Cauchy elasticity framework. Our analysis includes explicit solutions for both catenaries and longitudinal, piecewise constant stretched motions.
@@ 1. Introduction
This paper focuses on the physically important but simpler setting in which classical and modern advances in constitutive theory for elastic bodies can be explored: perfectly flexible strings. Interest in the study of the mechanics of strings and fibers has a storied history; it attracted the attention of da Vinci, Galileo, Euler, the Bernoullis, Savart, Lagrange, Leibniz, Ricatti, Huygens, Navier, Poisson, Lamé and other distinguished scientists. The interested reader can find a detailed account of the history of investigations into the mechanics of strings/fibers/wires in the work of Todhunter [36]. All of these early studies were within the context of special theories and not within the purview of general nonlinear theories of elasticity. Studies within the context of nonlinear theories are too numerous to mention and one can find copious references to both recent and older studies mentioned above in the treatise by Antman [1].
@@
In [31], the author established results on the existence and multiplicity of equilibria for stretch-limited catenaries, strings suspended between two points, a problem that attracted the interests of the likes of Galileo, Bernoulli, Huygens, Leibniz, among others. He found that the number of solutions for such stretch-limited strings deviates significantly from the case of traditional classical elastic strings satisfying (1.1). The work [32] studied the propagation of longitudinal waves in a semi-infinite stretch-limited string with the finite end held fixed and the other end asymptotically in tension. There, he proved the orbital asymptotic stability of an explicit two-parameter family of piecewise constant stretched motions.
@@
In this work, we introduce a simple class of stretch-limiting, piecewise linear constitutive relations for elastic strings of the form $\nu=\hat{\nu}(N)$. The organization of the paper is as follows. In Section 2, we briefly review the kinematics, balance laws and constitutive theory governing perfectly flexible elastic strings. In Section 3, we present a simple form of $\hat{\nu}(N)-1$ that limits the amount of compression and extension the string can undergo, is linear in $N$ when the string is extensible, and has different tangent moduli (i.e., proportionality constants) for extension versus compression (see (3.2)). In Section 4, we formulate simple boundary value problems, which are then solved from a semi-inverse perspective in Section 5 (for catenaries) and Section 6 (for piecewise constant stretched motions). Finally, in Section 7, we discuss potential applications of the proposed model.
The problems solved in Sections 5 and 6 illustrate key differences between the present work and previous studies while also offering potential applications. Section 5 provides explicit solutions for catenaries, unlike the more abstract results on the existence and multiplicity of solutions obtained in [31]. Similarly, the explicit piecewise constant stretch motions obtained in Section 6 involve a compressed segment and an elongated segment with distinct tangent moduli, whereas the solutions in [32] involve a maximally stretched segment and an elongated but still flexible segment. One application of the solutions in Section 5 is in modeling quasi-static deformations of stretch-limited strings, which may be relevant in civil engineering. In such cases, allowing for some elasticity while imposing a stretch limit offers insight into how these strings behave differently from the standard inextensible assumption often used in structural analysis (see, for example, Section 5/8 of [17]). The solutions in Sections 5 and 6 also serve as useful benchmarks for numerical methods, which are essential for studying more realistic problems involving these new constitutive theories. Future work will explore some of these numerical methods. In summary, the solutions obtained in this work offer new concrete perspectives on stretch-limited strings and open the door to further analytical and computational developments.
@@ 4.1. Alternative forms of the field equations
In the case that $N<0$, the only modification we need to consider to the above equation is to use $\hat{\nu}(-|\boldsymbol{n}(\xi,t)|)$ instead of $\hat{\nu}(|\boldsymbol{n}(\xi,t)|)$.
@@ 5. Catenaries
For each component ${r}_{i}$ of ${\boldsymbol{r}}$ (the components of ${\boldsymbol{n}}_{0}$ are ${n}_{0,i}$), we obtain
${r}_{1}({s})$ $=$ ${n}_{0,1}\ell({s}),\quad{r}_{3}({s})={n}_{0,3}\ell({s}),$ (5.6)
${r}_{2}({s})$ $=$ $\frac{1}{2\zeta}\left[2\sqrt{{n}_{0,1}^{2}+\left({n}_{0,2}+\zeta{s}\right)^{2}+{n}_{0,3}^{2}}+\zeta{s}\left(2{n}_{0,2}+\zeta{s}\right)-2\sqrt{{n}_{0,1}^{2}+{n}_{0,2}^{2}+{n}_{0,3}^{2}}\right],$ (5.7)
@@
$\ell({s})={s}+\frac{1}{\zeta}\ln\left[\frac{{n}_{0,2}+\zeta{s}+\sqrt{{n}_{0,1}^{2}+\left({n}_{0,2}+\zeta{s}\right)^{2}+{n}_{0,3}^{2}}}{{n}_{0,2}+\sqrt{{n}_{0,1}^{2}+{n}_{0,2}^{2}+{n}_{0,3}^{2}}}\right].$ (5.8)
@@
which can be used to find ${n}_{0,1}$, ${n}_{0,2}$ and ${n}_{0,3}$.
@@
In this section we study the case when there exist $\tilde{{s}}$ and $\breve{{s}}$ such that $10<\tilde{{s}}<1-\breve{{s}}<$ 1 and
@@
${r}_{1}({s})$ $=$ ${n}_{0,1}\ell_{A}({s}),\quad{r}_{3}({s})={n}_{0,3}\ell_{A}({s}),$ (5.14)
${r}_{2}({s})$ $=$ $\frac{\nu_{1}}{\zeta}\left[\sqrt{{n}_{0,1}^{2}+\left({n}_{0,2}+\zeta{s}\right)^{2}+{n}_{0,3}^{2}}-\sqrt{{n}_{0,1}^{2}+{n}_{0,2}^{2}+{n}_{0,3}^{2}}\right],$ (5.15)
@@
$\ell_{A}({s})=\frac{\nu_{1}}{\zeta}\ln\left[\frac{{n}_{0,2}+\zeta{s}+\sqrt{{n}_{0,1}^{2}+\left({n}_{0,2}+\zeta{s}\right)^{2}+{n}_{0,3}^{2}}}{{n}_{0,2}+\sqrt{{n}_{0,1}^{2}+{n}_{0,2}^{2}+{n}_{0,3}^{2}}}\right].$ (5.16)
@@
${r}_{1}({s})$ $=$ ${n}_{0,1}\ell_{B}({s})+{r}_{0,1},\quad{r}_{3}({s})={n}_{0,3}\ell_{B}({s})+{r}_{0,3},$ (5.18)
${r}_{2}({s})$ $=$ $\frac{1}{2\zeta}\left[2\sqrt{{n}_{0,1}^{2}+\left({n}_{0,2}+\zeta{s}\right)^{2}+{n}_{0,3}^{2}}+\zeta{s}\left(2{n}_{0,2}+\zeta{s}\right)\right.$ (5.19)
$\left.-2\sqrt{{n}_{0,1}^{2}+\left({n}_{0,2}+\zeta\tilde{{s}}\right)^{2}+{n}_{0,3}^{2}}-\zeta\tilde{{s}}\left(2{n}_{0,2}+\zeta\tilde{{s}}\right)\right]+{r}_{0,2},$
@@
$\ell_{B}({s})={s}-\tilde{{s}}+\frac{1}{\zeta}\ln\left[\frac{{n}_{0,2}+\zeta{s}+\sqrt{{n}_{0,1}^{2}+\left({n}_{0,2}+\zeta{s}\right)^{2}+{n}_{0,3}^{2}}}{{n}_{0,2}+\zeta\tilde{{s}}+\sqrt{{n}_{0,1}^{2}+\left({n}_{0,2}+\zeta\tilde{{s}}\right)^{2}+{n}_{0,3}^{2}}}\right].$ (5.20)
@@
${r}_{1}({s})$ $=$ ${n}_{0,1}\ell_{C}({s})+\hat{{r}}_{0,1},\quad{r}_{3}({s})={n}_{0,3}\ell_{C}({s})+\hat{{r}}_{0,3},$ (5.22)
${r}_{2}({s})$ $=$ $\frac{\nu_{1}}{\zeta}\left\{\sqrt{{n}_{0,1}^{2}+\left({n}_{0,2}+\zeta{s}\right)^{2}+{n}_{0,3}^{2}}-\sqrt{{n}_{0,1}^{2}+\left[{n}_{0,2}+\zeta\left(1-\breve{{s}}\right)\right]^{2}+{n}_{0,3}^{2}}\right\}+\hat{{r}}_{0,2},$ (5.23)
@@
$\ell_{C}({s})=\frac{\nu_{1}}{\zeta}\ln\left\{\frac{{n}_{0,2}+\zeta{s}+\sqrt{{n}_{0,1}^{2}+\left({n}_{0,2}+\zeta{s}\right)^{2}+{n}_{0,3}^{2}}}{{n}_{0,2}+\zeta{s}+\sqrt{{n}_{0,1}^{2}+\left[{n}_{0,2}+\zeta\left(1-\breve{{s}}\right)\right]^{2}+{n}_{0,3}^{2}}}\right\}.$ (5.24)
In the above results for ${r}_{i}({s})$ we have the following list of 11 unknown parameters: $\tilde{{s}}$, $\breve{{s}}$, ${r}_{0,1}$, ${r}_{0,2}$, ${r}_{0,3}$, $\hat{{r}}_{0,1}$, $\hat{{r}}_{0,2}$, $\hat{{r}}_{0,3}$, ${n}_{0,1}$, ${n}_{0,2}$ and ${n}_{0,3}$. These can be found from the 6 continuity conditions
@@
${n}_{0,1}^{2}+\left({n}_{0,2}+\zeta\tilde{{s}}\right)^{2}+{n}_{0,3}^{2}={N}_{1}^{2},\quad{n}_{0,1}^{2}+\left[{n}_{0,2}+\zeta\left(1-\breve{{s}}\right)\right]^{2}+{n}_{0,3}^{2}={N}_{1}^{2}.$ (5.27)
From (5.25), (5.26), and (5.27), we have 11 equations determining the list of 11 unknowns.
@@ 7. Final remarks
In the present work, we proposed a new type of piecewise linear constitutive relation for strings that limits the amount of extension and compression the string can undergo. Moreover, the tangent moduli relating tension to stretch are different for extension versus compression. We solved static and dynamic problems from the semi-inverse perspective with solutions to the latter having piecewise constant stretch. Numerical methods for solving more realistic static problems will be explored in future work, while dynamic problems will be studied using the method developed in [3].
The constitutive model developed in this paper provides a simple framework for describing the behavior of stretch-limited strings, which appear in various physical and biological settings and cannot be modeled using the classical approach. As shown in Figure 1 of [21], in general, implicit constitutive theories are particularly useful in situations where the conventional approach of expressing the stress (or, in this case, the tension) explicitly as a function of the strain is inadequate. This is precisely the case for the constitutive relation (3.2) introduced here: when the stretch reaches a critical threshold, it remains fixed, yet the tension can continue to either decrease or increase without bound. This behavior fundamentally defies modeling within the classical framework, which expresses tension as a function of stretch. The model developed in this study has potential applications in biological systems, where stretch-limited strings arise in different contexts (see, e.g., [13], particularly Figures 1–3, and the references therein). More broadly, the model presented here offers a simple alternative to classical elastic string models where standard constitutive assumptions fail to incorporate natural stretch-limiting behavior.